# Movable Antenna Enhanced Networked Full-Duplex Integrated Sensing and Communication System

### v2

@@ III-I Optimizing DL User MA’s Position
[…]
@@
Obviously, obtaining the optimal solution to problem (P13) is challenging due to the non-convex nature of the objective function (78a). Therefore, we still employ the MM methodology to address it.
Firstly, the term $(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{1}_{k,m,n}\mathbf{h}^{0,0}_{n,k}$ in (78a) can be upper-bounded by invoking (63) in Lemma 1 as shown in (80), where $c_{6,2,k,m,n,i,j}$ is constant, $\mathbf{q}^{3}_{k,m,n,i,j}\triangleq\nabla h^{R1}_{k,m,n,i,j}-\psi_{1,k,m,n,i,j}\mathbf{r}_{0,k,0}$, $\mathbf{r}_{0,k,0}=[x^{d}_{k,0},y^{d}_{k,0}]^{T}$ is obtained form the last iteration, $\nabla h^{R1}_{k,m,n,i,j}$ is the the gradient vector of $h^{R1}_{k,m,n,i,j}$ over $\mathbf{r}_{0,k,0}$, $\psi_{1,k,m,n,i,j}=\frac{32\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{1}_{k,m,n}[i,j]|$, and the derivations of $\nabla h^{R1}_{k,m,n,i,j}$ and $\psi_{1,k,m,n,i,j}$ are presented in Appendix A.
$(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{4}_{k,m,n}\mathbf{h}^{0,0}_{n,k}={\sum}_{i=1}^{L^{r,0}_{m,k}}{\sum}_{j=1}^{L^{r,0}_{n,k}}(\mathbf{h}^{0,0}_{m,k}[i])^{\ast}\mathbf{Q}^{4}_{k,m,n}[i,j]\mathbf{h}^{0,0}_{m,k}[j]$ (80)
$={\sum}_{i=1}^{L^{r,0}_{m,k}}{\sum}_{j=1}^{L^{r,0}_{n,k}}\underbrace{|\mathbf{Q}^{4}_{k,m,n}[i,j]|\cos(\frac{2\pi}{\lambda}(\mathbf{r}_{0,k})^{T}(-\mathbf{a}^{r,0}_{m,k,i}+\mathbf{a}^{r,0}_{n,k,j})+\angle\mathbf{Q}^{4}_{k,m,n}[i,j])}\limits_{h^{R1}_{k,m,n,i,j}}$
$\overset{(a)}{\leq}{\sum}_{i=1}^{L^{r,0}_{m,k}}{\sum}_{j=1}^{L^{r,0}_{n,k}}\frac{\psi_{1,k,m,n,i,j}}{2}\|\mathbf{r}_{0,k}\|^{2}_{2}+(\mathbf{q}^{3}_{k,m,n,i,j})^{T}\mathbf{r}_{0,k}+c_{6,2,k,m,n,i,j},$
Furthermore, the terms $(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{2}_{k,m}\mathbf{h}^{0,0}_{m,k}-\textrm{ Re }\{(\mathbf{q}^{1}_{k,m})^{H}\mathbf{h}^{0,0}_{m,k}\}$ and $(\mathbf{h}^{3}_{k,l})^{H}\mathbf{Q}^{3}_{k,l}\mathbf{h}^{3}_{k,l}$ all can be similarly convexified following the same arguments as in (61) and (64) in Sec III-G. Therefore, the upper-bounds of $(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{2}_{k,m}\mathbf{h}^{0,0}_{m,k}-\textrm{ Re }\{(\mathbf{q}^{1}_{k,m})^{H}\mathbf{h}^{0,0}_{m,k}\}$ and $(\mathbf{h}^{3}_{k,l})^{H}\mathbf{Q}^{3}_{k,l}\mathbf{h}^{3}_{k,l}$ are shown in (81) and (82), respectively, in the following
$(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{2}_{k,m}\mathbf{h}^{0,0}_{m,k}-\textrm{ Re }\{(\mathbf{q}^{1}_{k,m})^{H}\mathbf{h}^{0,0}_{m,k}\}$ (81)
$\leq\frac{\psi_{2,k,m}}{2}\|\mathbf{r}_{0,k}\|^{2}_{2}+(\mathbf{q}^{3}_{k,m})^{T}\mathbf{r}_{0,k}+c_{6,2,k,m},$
$(\mathbf{h}^{3}_{k,l})^{H}\mathbf{Q}^{3}_{k,l}\mathbf{h}^{3}_{k,l}\leq\frac{\psi_{3,k,l}}{2}\|\mathbf{r}_{0,k}\|^{2}_{2}+(\mathbf{q}^{4}_{k,l})^{T}\mathbf{r}_{0,k}+c_{6,3,k,l},$ (82)
with $c_{6,2,k,m}$ and $c_{6,3,k,l}$ being constants, the numbers $\psi_{2,k,m}$ and $\psi_{3,k,l}$ adhere to the condition outlined in Lemma 1.
In accordance with the aforementioned MM transformation, the update of $\mathbf{r}_{0,k}$ is achieved by solving a convex surrogate of the original problem (P13), as follows
[…]
where $c_{6,4,k}$ is constant, $\psi_{4,k}\!=\!0.5({\sum}_{m=1}^{M_{t}}\psi_{2,k,m}\!+\!\!{\sum}_{m=1}^{M_{t}}{\sum}_{n=1}^{M_{t}}{\sum}_{i=1}^{L^{r,0}_{m,k}}{\sum}_{j=1}^{L^{r,0}_{n,k}}\psi_{1,k,m,n,i,j}\!+\!\!{\sum}_{l=1}^{K_{u}}\psi_{3,k,l})$ and $\mathbf{q}^{5}_{k}\!=\!{\sum}_{m=1}^{M_{t}}{\sum}_{n=1}^{M_{t}}{\sum}_{i=1}^{L^{r,0}_{m,k}}{\sum}_{j=1}^{L^{r,0}_{n,k}}\mathbf{q}^{2}_{k,m,n,i,j}\!+\!\!{\sum}_{m=1}^{M_{t}}\mathbf{q}^{3}_{k,m}\!+\!\!{\sum}_{l=1}^{K_{u}}\mathbf{q}^{4}_{k,l}$. The problem (P14) is convex and can be solved by CVX.
Since the units of $\mathbf{r}_{0,k}$, i.e., $x^{d}_{k}$ and $y^{d}_{k}$, are not coupled, the problem w.r.t. $x^{d}_{k}$ can be written as
[…]
@@
Since $\psi_{4,k}>0$, we can obtain the optimal solution $(x^{d}_{k})^{\star}$ by directly determining the position of the axis of symmetry of the function (84a), which are given as
$(x^{d}_{k})^{\star}\!=\!\left\{\begin{aligned} -{A}/{2},\ &\textrm{if}\ -{\mathbf{q}^{5}_{k}[1]}/({2\psi_{4,k}})<-{A}/{2},\\ -{\mathbf{q}^{5}_{k}[1]}/{2\psi_{4,k}},\ &\textrm{if}\ -{A}/{2}\leq-{\mathbf{q}^{5}_{k}[1]}/{2\psi_{4,k}}\leq{A}/{2},\\ {A}/{2},\ &\textrm{if}\ -{A}/{2}<-{\mathbf{q}^{5}_{k}[1]}/{2\psi_{4,k}}.\end{aligned}\right.$ (85)
@@
$(y^{d}_{k})^{\star}\!=\!\left\{\begin{aligned} -{A}/{2},\ &\textrm{if}\ -{\mathbf{q}^{5}_{k}[2]}/({2\psi_{4,k}})<-{A}/{2},\\ -{\mathbf{q}^{5}_{k}[1]}/{2\psi_{4,k}},\ &\textrm{if}\ -{A}/{2}\leq-{\mathbf{q}^{5}_{k}[2]}/{2\psi_{4,k}}\leq{A}/{2},\\ {A}/{2},\ &\textrm{if}\ -{A}/{2}<-{\mathbf{q}^{5}_{k}[2]}/{2\psi_{4,k}}.\end{aligned}\right.$ (86)
@@ III-J Optimizing UL User MA’s Position
When other variables are obtained, the optimization problem solving for the MA’s position ${\mathbf{r}_{1,l}}$ of the $l$-th UL user is formulated as
[…]
$+{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{5}_{l,m,n}\mathbf{h}^{0,1}_{n,l}+c_{7,1,l}$ (87a)
s.t. ${\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{6}_{l,m,n}\mathbf{h}^{0,1}_{n,l}\!+\!c_{7,2,l}\leq 0,$ (87b)
@@
with $c_{7,1,l}$ and $c_{7,2,l}$ being constant terms, and we define
$\mathbf{q}^{6}_{l,m}\triangleq 2\omega_{d,k}\beta_{d,k}\mathbf{\Sigma}^{1}_{m,l}\mathbf{H}^{1}_{m,i}\mathbf{u}_{m,l},$ (88)
$\mathbf{Q}^{6}_{l,m,n}\triangleq\mathbf{\Sigma}^{1}_{m,i}\mathbf{H}^{1}_{m,i}\mathbf{u}_{m,0}\mathbf{u}_{n,0}^{H}(\mathbf{H}^{1}_{n,i})^{H}(\mathbf{\Sigma}^{1}_{n,i})^{H},$
Unfortunately, this problem is still difficult to solve due to the non-convex objective (87a) and constraint (87b) w.r.t $\mathbf{r}_{1,l}$. To address these issues, we will construct the convex surrogate functions of (87a) and (87b) by applying MM method, and then recast the problem (P16) into an explicit form of $\mathbf{r}_{1,l}$.
Firstly, by using the same MM-based algorithmic framework in (80), the convex upper-bounds of the terms $(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{5}_{l,m,n}\mathbf{h}^{0,1}_{n,l}$ and $(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{6}_{l,m,n}\mathbf{h}^{0,1}_{n,l}$ are respectively provided as
$(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{5}_{l,m,n}\mathbf{h}^{0,1}_{n,l}\leq{\sum}_{i=1}^{L^{r,1}_{m,l}}{\sum}_{j=1}^{L^{r,1}_{n,l}}\frac{\psi_{6,l,m,n,i,j}}{2}\|\mathbf{r}_{1,l}\|^{2}_{2}$ (89)
$+(\mathbf{q}^{8}_{l,m,n,i,j})^{T}\mathbf{r}_{1,l}+c_{7,3,l,m,n,i,j},$
$(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{6}_{l,m,n}\mathbf{h}^{0,1}_{n,l}\leq{\sum}_{i=1}^{L^{r,1}_{m,l}}{\sum}_{j=1}^{L^{r,1}_{n,l}}\frac{\psi_{9,l,m,n,i,j}}{2}\|\mathbf{r}_{1,l}\|^{2}_{2}$ (90)
$+(\mathbf{q}^{10}_{l,m,n,i,j})^{T}\mathbf{r}_{1,l}+c_{7,4,l,m,n,i,j},$
with $c_{7,3,l,m,n,i,j}$ and $c_{7,4,l,m,n,i,j}$ being constant, $\psi_{6,l,m,n,i,j}=\frac{32\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{10}_{l,m,n}[i,j]|$ and $\psi_{9,l,m,n,i,j}=\frac{32\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{11}_{l,m,n}[i,j]|$.
Furthermore, we can construct an upper-bound for $(\mathbf{h}^{2}_{k,l})^{H}\mathbf{Q}^{4}_{k,l}\mathbf{h}^{2}_{k,l}$ using the same operations as in (61) and (64), as follows
$(\mathbf{h}^{2}_{k,l})^{H}\mathbf{Q}^{4}_{k,l}\mathbf{h}^{2}_{k,l}\leq\frac{\psi_{5,k,l}}{2}\|\mathbf{r}_{1,l}\|^{2}_{2}+(\mathbf{q}^{7}_{k,l})^{T}\mathbf{r}_{1,l}+c_{7,5,k,l},$ (91)
where $c_{7,5,k,l}$ is constant, and $\psi_{5,k,l}$ and $\mathbf{q}^{7}_{k,l}$ are derived using similar arguments as presented in in (61) and (64).
Subsequently, referring to (68), a convex surrogate function for the term $\textrm{Re}\{\mathbf{q}^{6}_{l,m}\mathbf{h}^{0,1}_{m,l}\}$ can be derived as follows
$\textrm{Re}\{\mathbf{q}^{6}_{l,m}\mathbf{h}^{0,1}_{m,l}\}\!\geq\!-\frac{\psi_{7,l,m}}{2}\|\mathbf{r}_{1,l}\|^{2}_{2}+(\mathbf{q}^{9}_{l,m})^{T}\mathbf{r}_{1,l}+c_{7,6,l,m},$ (92)
where $c_{7,6,l,m}$ denotes a constant term, and $\psi_{7,l,m}$ and $\mathbf{q}^{9}_{l,m}$ can be obtained leveraging the results presented in (68).
Substituting the result in (89)-(92) into the original problem (P16), we further obtain
[…]
s.t. $\psi_{10,l}\|\mathbf{r}_{1,l}\|^{2}_{2}+(\mathbf{q}^{12}_{l})^{T}\mathbf{r}_{1,l}+c_{7,8,l}\leq 0,$ (93b)
$\mathbf{r}_{1,l}\in\mathcal{C},$ (93c)
where $c_{7,7,l}$ and $c_{7,8,l}$ are constants, $\psi_{8,l}\triangleq{\sum}_{k=1}^{K_{d}}\frac{\psi_{5,k,l}}{2}+{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}{\sum}_{i=1}^{L^{r,1}_{m,l}}{\sum}_{j=1}^{L^{r,1}_{n,l}}\frac{\psi_{6,l,m,n,i,j}}{2}+{\sum}_{m=1}^{M_{r}}\frac{\psi_{7,l,m}}{2}$, $\mathbf{q}^{10}_{l}\triangleq{\sum}_{k=1}^{K_{d}}\mathbf{q}^{7}_{k,l}+{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}{\sum}_{i=1}^{L^{r,1}_{m,l}}{\sum}_{j=1}^{L^{r,1}_{n,l}}\mathbf{q}^{8}_{l,m,n,i,j}-{\sum}_{m=1}^{M_{r}}\mathbf{q}^{9}_{l,m}$, $\psi_{10,l}\triangleq{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}{\sum}_{i=1}^{L^{r,1}_{m,l}}{\sum}_{j=1}^{L^{r,1}_{n,l}}\frac{\psi_{9,l,m,n,i,j}}{2}$ and $\mathbf{q}^{12}_{l}\!\!\triangleq\!\!{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}{\sum}_{i=1}^{L^{r,1}_{m,l}}{\sum}_{j=1}^{L^{r,1}_{n,l}}\mathbf{q}^{11}_{l,m,n,i,j}$. Now, the problem (P17) is convex that can be solved optimally by CVX.
Moreover, the overall algorithm to solve (P1) is specified in Algorithm 1.
[…]
@@
2:  randomly generate feasible $\{\mathbf{w}^{0}_{m,k},(\mathbf{W}^{r}_{m})^{0}\}$, $\{\mathbf{u}_{p,j}^{0}\}$$\{q_{l}\}$$\{(\mathbf{t}^{0}_{m,n})^{0},(\mathbf{t}^{1}_{m,n})^{0}\}$ and $\{\mathbf{r}_{0,k}^{0},\mathbf{r}_{1,l}^{0}\}$;
@@
13:  until $convergence$

### v3

@@ III-I Optimizing DL User MA’s Position
$\textrm{(P13)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{r}^{0}_{k}\in\mathcal{C}}{\sum}_{m=1}^{M_{t}}{\sum}_{n=1}^{M_{t}}(\mathbf{h}^{0,0}_{m,k})^{H}\mathbf{Q}^{1}_{k,m,n}\mathbf{h}^{0,0}_{n,k}$ (94a)
@@
Obviously, obtaining the optimal solution to the problem (P13) is challenging due to the non-convex nature of the objective function (94a). Therefore, we still employ the MM methodology to convexify it. Due to space constraints, we omit the transformation process.
The update of $\mathbf{r}_{0,k}$ is achieved by solving a convex surrogate of the original problem (P13), as follows
$\textrm{(P14)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{r}_{0,k}\in\mathcal{C}}{\psi_{7,k}}\|\mathbf{r}_{0,k}\|^{2}_{2}+(\mathbf{q}^{2}_{k})^{T}\mathbf{r}_{0,k}+c_{6,2,k}$ (96a)
The problem (P14) is convex and can be solved by CVX.
Since the units of $\mathbf{r}_{0,k}$, i.e., $x^{d}_{k}$ and $y^{d}_{k}$, are not coupled, the subproblem w.r.t. $x^{d}_{k}$ can be written as
$\textrm{(P15)}:$ $\mathop{\textrm{min}}\limits_{x^{d}_{k}}{\psi_{7,k}}(x^{d}_{k})^{2}+\mathbf{q}^{2}_{k}[1]x^{d}_{k}$ (97a)
@@
Since $\psi_{7,k}>0$, we can obtain the optimal solution $(x^{d}_{k})^{\star}$ by directly determining the position of the axis of symmetry of the function (97a), which are given as
$(x^{d}_{k})^{\star}\!=\!\left\{\begin{aligned} -{A}/{2},\ &\textrm{if}\ -{\mathbf{q}^{2}_{k}[1]}/({2\psi_{7,k}})<-{A}/{2},\\ -{\mathbf{q}^{2}_{k}[1]}/({2\psi_{7,k}}),\ &\textrm{if}\ -{A}/{2}\leq-{\mathbf{q}^{2}_{k}[1]}/({2\psi_{7,k}})\leq{A}/{2},\\ {A}/{2},\ &\textrm{if}\ -{A}/{2}<-{\mathbf{q}^{2}_{k}[1]}/({2\psi_{7,k}}).\end{aligned}\right$. (98)
@@
$(y^{d}_{k})^{\star}\!=\!\left\{\begin{aligned} -{A}/{2},\ &\textrm{if}\ -{\mathbf{q}^{2}_{k}[2]}/({2\psi_{7,k}})<-{A}/{2},\\ -{\mathbf{q}^{2}_{k}[1]}/({2\psi_{7,k}}),\ &\textrm{if}\ -{A}/{2}\leq-{\mathbf{q}^{2}_{k}[2]}/({2\psi_{7,k}})\leq{A}/{2},\\ {A}/{2},\ &\textrm{if}\ -{A}/{2}<-{\mathbf{q}^{2}_{k}[2]}/({2\psi_{7,k}}).\end{aligned}\right$. (99)
@@ III-J Optimizing UL User MA’s Position
In this subsection, we will investigate the update of the MA’s position $\mathbf{r}_{1,l}$} for the $l$-th UL user. Firstly, the functions $\tilde{\mathrm{R}}_{d,k}$ and $\tilde{\mathrm{R}}_{u,l}$ can be respectively rewritten as
${\sum}_{k=1}^{K_{d}}\tilde{\mathrm{R}}_{d,k}\!\triangleq\!{\sum}_{k=1}^{K_{d}}\!-\!(\mathbf{h}^{2}_{k,l}(\mathbf{r}_{1,l}))^{H}\mathbf{Q}^{4}_{k,l}\mathbf{h}^{2}_{k,l}(\mathbf{r}_{1,l})\!+\!c_{7,1,l},$ (100)
${\sum}_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}\!\triangleq\!\!{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}(\mathbf{h}^{0,1}_{m,l}(\mathbf{r}_{1,l}))^{H}\mathbf{Q}^{5}_{l,m,n}\mathbf{h}^{0,1}_{n,l}(\mathbf{r}_{1,l})$ (101)
$-{\sum}_{m=1}^{M_{r}}\textrm{Re}\{\mathbf{q}^{3}_{l,m}\mathbf{h}^{0,1}_{m,l}(\mathbf{r}_{1,l})\}+c_{7,2,l},$
with $c_{7,1,l}$ and $c_{7,2,l}$ being constant terms, and we define
$\mathbf{q}^{3}_{l,m}\triangleq 2\omega_{d,k}\beta_{d,k}\mathbf{\Sigma}^{1}_{m,l}\mathbf{H}^{1}_{m,i}\mathbf{u}_{m,l},$ (102)
Furthermore, the sensing SINR constraint (30b) can be equivalent rewritten as
$\textrm{SINR}_{t}\geq\Gamma_{r}\Longleftrightarrow$ (103)
${\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}(\mathbf{h}^{0,1}_{m,l}(\mathbf{r}_{1,l}))^{H}\mathbf{Q}^{6}_{l,m,n}\mathbf{h}^{0,1}_{n,l}(\mathbf{r}_{1,l})+c_{7,3,l},$
where $c_{7,3,l}$ is constant and $\mathbf{Q}^{6}_{l,m,n}\triangleq\mathbf{\Sigma}^{1}_{m,l}\mathbf{H}^{1}_{m,l}\mathbf{u}_{m,0}\mathbf{u}_{n,0}^{H}(\mathbf{H}^{1}_{n,l})^{H}(\mathbf{\Sigma}^{1}_{n,l})^{H}$.
Therefore, the problem w.r.t. $\mathbf{r}_{1,l}$ can be rewritten as
$\textrm{(P16)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{r}_{1,l}}\!{\sum}_{k=1}^{K_{d}}\!(\mathbf{h}^{2}_{k,l})^{H}\mathbf{Q}^{4}_{k,l}\mathbf{h}^{2}_{k,l}\!\!-\!\!\!{\sum}_{m=1}^{M_{r}}\textrm{Re}\{\mathbf{q}^{3}_{l,m}\mathbf{h}^{0,1}_{m,l}\}$
$+{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{5}_{l,m,n}\mathbf{h}^{0,1}_{n,l}+\hat{c}_{7,1,l}$ (104a)
s.t. $\!{\sum}_{m=1}^{M_{r}}{\sum}_{n=1}^{M_{r}}(\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{Q}^{6}_{l,m,n}\mathbf{h}^{0,1}_{n,l}\!+\!c_{7,3,l}\!\leq\!0,$ (104b)
@@
with $\hat{c}_{7,1,l}$ being a constant term.
Unfortunately, this problem is still difficult to solve due to the non-convex objective (104a) and constraint (104b) w.r.t $\mathbf{r}_{1,l}$. To address these issues, we will construct the convex surrogate functions of (104a) and (104b) by applying the MM method, which the derivation details are omitted for the space constraints, and then recast the problem (P11) into an explicit form of $\mathbf{r}_{1,l}$, which is given as
$\textrm{(P17)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{r}_{1,l}}\ \psi_{8,l}\|\mathbf{r}_{1,l}\|^{2}_{2}+(\mathbf{q}^{4}_{l})^{T}\mathbf{r}_{1,l}+c_{7,4,l}$ (105a)
s.t. $\psi_{10,l}\|\mathbf{r}_{1,l}\|^{2}_{2}+(\mathbf{q}^{5}_{l})^{T}\mathbf{r}_{1,l}+c_{7,5,l}\leq 0,$ (105b)
$\mathbf{r}_{1,l}\in\mathcal{C}.$ (105c)
Now, the problem (P17) is convex, which can be solved optimally by CVX.
Moreover, the overall algorithm to solve (P1) is specified in Algorithm 1. $\eta=\{\{\mathbf{w}_{m,k},\mathbf{W}^{r}_{m}\},\{\mathbf{u}_{p,i}\},\{q_{l}\},\{\mathbf{t}^{0}_{m,n},\mathbf{t}^{1}_{p,j},\mathbf{r}_{0,k},\mathbf{r}_{1,l}\}\}$ denote the collection of all variables. And let $g(\eta)$ represent the sum-rate for both downlink and uplink users.
Algorithm 1 Overall Algorithm to Solve $(\mathrm{P1})$
@@
2: randomly generate feasible $\{\mathbf{w}^{0}_{m,k},(\mathbf{W}^{r}_{m})^{0}\}$, $\{\mathbf{u}_{p,j}^{0}\}$, $\{q_{l}\}$, $\{(\mathbf{t}^{0}_{m,n})^{0},(\mathbf{t}^{1}_{m,n})^{0}\}$ and $\{\mathbf{r}_{0,k}^{0},\mathbf{r}_{1,l}^{0}\}$;
@@
13:until $\bigg|\frac{g(\eta^{i+1})-g(\eta^{i})}{g(\eta^{i})}\bigg|\leq\epsilon,$
III-K Complexity
In the following, we will discuss the complexity of our proposed algorithms. According to the complexity analysis in [45], the complexity of solving SOCP problems (P3), (P10), (P12) and (P17) are $\mathcal{O}(C_{1}M_{t}^{3.5}N_{t}^{6})$, $\mathcal{O}(C_{2}M_{t}N_{t}^{1.5})$, $\mathcal{O}(C_{3}M_{r}N_{r}^{1.5})$ and $\mathcal{O}(76C_{4})$, respectively, where $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$ are respectively the number of iterations of solving problem (P3), (P10), (P12) and (P17) by the MM method. The complexity of solving (P4) is $\mathcal{O}(K_{u}^{3.5})$. Therefore, the total computational complexity of Algorithm 1 is approximately given as $\mathcal{O}(C_{5}(C_{1}M_{t}^{3.5}N_{t}^{6}+K_{u}^{3.5}))$ with $C_{5}$ represented as the number of iterations to solve problem (P1).